OpenStudy (anonymous):

Solve the system for x. 9x + 2y = 11 -22x - 5y = -11 A. -143 B. -33 C. 33 D. 143

OpenStudy (anonymous):

9x + 2y = 11 *5 -22x - 5y = -11 *2 45x + 10y = 55 -44x - 10y = -22 by addition x=33